OpenStudy (anonymous):

use factorization to solve the quadratic equation 4x^2=36x-81

OpenStudy (anonymous):

4x^2-36x+81=0 => (2x-9)^2=0 =>2x=9=>x=9/2

OpenStudy (anonymous):

thanks!